A particle is moving in the +X direction with its velocity  V= .Assume that at t=0,the particle was located at X=0 then the value of time dependence of velocity

Grade:11

1 Answers

deepak
7 years ago
clearly state the value of v
i think i know the question,
if v=ax^1/2
then take v=dx/dt
=>dx/dt=ax^12
=>dx/(x^1/2)=at
on intergrating,
2x^1/2=at+c
as its give x=0 => t=0, substituting
c=0
hence,
2x^1/2=at
=>x^1/2=at/2
substituting this in the first eqn,
v=a2t/2
 hope this helps XP